Consider the demand for a good. At price Rs 4, the demand for the good is 25 units.
Suppose price of the good increases to Rs 5, and as a result, the demand for the good
falls to 20 units. Calculate the price elasticity?
1
Expert's answer
2021-10-05T11:18:58-0400

The price elasticity of demand is calculated as follows;


"(E \nd\n\u200b\n ) =(-) \\frac {P} {Q} \u00d7\\frac {\\Delta Q} {\\Delta P}"


Given, "P=Rs.4;P1\ufeff=Rs.5"


"\u0394P=P \n1\n\u200b\n \u2212P=Rs.5\u2212Rs.4=Rs.1"


"Q=25\\space units ;Q_{1} \ufeff=20units"


"\u0394Q=Q \n1\n\u200b\n \u2212Q=(20\u221225)\\space units=(\u2212)5\\space units"


"E \nd\n\u200b\n =(\u2212)\\frac {4}{25}\n \n\n\u200b\n \u00d7\\frac { -5}{1}"


"=-0.8"


Here the elasticity of demand smaller than 1 implies that demand is less elastic and the negative sign denotes the inverse relationship between quantity demanded and price of the commodity.